# Kudla Millson lift of toric cycles and restriction of Hilbert modular forms

### v3

@@ Acknowledgments.
for $x\in E$. Explicitely we have $\epsilon_{Q}(x)=A^{-1}_{Q}x^{T}A_{Q}$ where $A_{Q}$ denotes the Gram matrix of $Q$ and $x^{T}$ is the transpose of $x$. Let $E_{T}\subset\operatorname{End}(V)$ be the subalgebra consisting of all the $\mathbb{Q}$-endomorphisms $x\in\operatorname{End}(V)$ such that $xt=tx$ for any $t\in{\operatorname{T}}(\mathbb{Q})$. We will denote it simply by $E$ since there is no risk of confusion, and let $F$ be the subalgebra fixed by $\epsilon_{Q}$. We view $\operatorname{T}(\mathbb{Q})\subset\operatorname{SO}_{V}(\mathbb{Q})\subset\operatorname{End}(V)$. The following proposition is proved in [BCKM03, Proposition. 3.3].
@@
Over $\overline{\mathbb{Q}}$, the quadratic space $V$ is isomorphic to ¯$\overline{\mathbb{Q}}^{2d}$ with quadratic form $Q(v,v)=v_{1}v_{d+1}+v_{2}v_{d+2}+\cdots+v_{d}v_{2d}$ i.e. with Gram matrix $\begin{pmatrix}0&\mathbf{1}_{d}\\ \mathbf{1}_{d}&0\end{pmatrix}$. With respect to this quadratic form, the torus can be diagonalized in $\operatorname{SO}_{V}(\overline{\mathbb{Q}})$ to
@@ 3.1.  Weil representation
    the field $\mathcal{k}=F$ is an arbitrary totally real field of degree $k=d$. The quadratic space is $\mathcal{V}=E$, where $E$ is an étale algebra viewed as quadratic space of dimension $22m=$ 2 over $F$, and equipped with the quadratic form $Q_{\alpha}$. The orthogonal group $\operatorname{SO}_{\mathcal{V}}=\operatorname{SO}_{E}\simeq E^{1}$ is a torus.
@@
Let us describe this action more precisely, on a Schwartz function $\varphi=\varphi_{\infty}\otimes\varphi_{f}\in\mathcal{S}(\mathcal{V}_{\mathbb{R}})\otimes\mathcal{S}(\mathcal{V}_{\widehat{\mathbb{Q}}})$. For $h\in\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k}})$, we have
@@
$(\omega(g,1)\varphi)(v)$ $=\chi_{w}(b\mathcal{Q}(v,v))\varphi(v).$ (3.22)
@@
For every real place $\sigma$ of $\mathcal{k}$, let $\mathcal{V}_{\sigma}\coloneqq\mathcal{V}\otimes_{\sigma}\mathbb{R}$ and $\mathcal{V}_{\mathbb{R}}\coloneqq\bigoplus\mathcal{V}_{\sigma}$. It is a real quadratic space, and let $(p_{\sigma},q_{\sigma})$ be its signature. Let $\mathcal{D}_{\sigma}^{+}$ be the associated connected symmetric space, that can be described as one of the two connected of the Grassmanian $\mathcal{D}_{\sigma}$ of $q_{\sigma}$-dimensional oriented subspaces $z\subset\mathcal{V}_{\sigma}$ that are negative, i.e. ${\left.\kern-1.2pt\mathcal{Q}\vphantom{\big{|}}\right|_{z}}<0$. We set $\mathcal{D}^{+}=\prod\mathcal{D}_{\sigma}^{+}$, where the product ranges over the archimedean places of $\mathcal{k}$ for which $p_{\sigma}q_{\sigma}$ is nonzero [^2] In the case where $p_{\sigma}=0$ or $q_{\sigma}=0$, the manifold $\mathcal{D}_{\sigma}^{+}$ is just a point.. The dimension of $\mathcal{D}^{+}_{\sigma}$ is $p_{\sigma}q_{\sigma}$, so that the dimension of $\mathcal{D}^{+}$ is $\sum_{\sigma}p_{\sigma}q_{\sigma}$. At every place $\sigma$ let $\mathcal{k}_{\sigma}$ be the completion of $\mathcal{k}$, so that $\mathcal{k}_{\infty}\simeq\prod_{\sigma}\mathcal{k}_{\sigma}$. Since we assumed that $\mathcal{k}$ only has real places, we have $\mathcal{k}_{\sigma}\simeq\mathbb{R}$. At the level of the orthogonal group, we have $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\infty})\simeq\prod_{\simeq}\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\sigma})$, where $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\sigma})\simeq\operatorname{SO}(p_{\sigma},q_{\sigma})$. The connected component $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\infty})^{+}\simeq\prod_{\sigma}\operatorname{SO}(p_{\sigma},q_{\sigma})^{+}$ of the identity acts transitively on $\mathcal{D}^{+}$. Let $K_{\infty}$ be a maximal connected compact subgroup of $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\infty})^{+}$, that is isomorphic to $\prod_{\sigma}\operatorname{SO}(p_{\sigma})\times\operatorname{SO}(q_{\sigma})$. Hence
@@
is the space mentionned in the introduction. It has two connected components (corresponding to the two choices of orientations) and $\mathcal{D}^{+}$ is one of them.
@@
Let $\varphi_{f}\in\mathcal{S}(\mathcal{V}_{\widehat{\mathcal{k}}})$ be a finite Schwartz function, let $K_{f}\subset\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{O}_{\mathcal{k}}})$ be an open compact preserving $\varphi_{f}$ in the sense that $\omega(k)\varphi_{f}=\varphi_{f}$. Let $K\coloneqq K_{\infty}K_{f}$ and consider the adelic space
$Y=\operatorname{SO}_{\mathcal{V}}(\mathcal{k})\backslash\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k}})/K\simeq\operatorname{SO}_{\mathcal{V}}(\mathcal{k})\backslash\mathcal{D}\times\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}})/K_{f}.$ (3.27)
@@
$\mathcal{C}\coloneqq\operatorname{SO}_{\mathcal{V}}(\mathcal{k})^{+}\backslash\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}})/K_{f}.$ (3.28)
@@
$\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}})=\bigsqcup_{I\in\mathcal{C}}\operatorname{SO}_{\mathcal{V}}(\mathcal{k})^{+}h_{I}K_{f}.$ (3.29)
where $h_{I}\in\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}})$ are representants. The adelic space is a finite union of locally symmetric spaces
@@
Let $C^{\infty}(\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}}))$ be the space of locally constant functions on $\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}})$. The map
$\left[\Omega^{r}(\mathcal{D}^{+})\otimes_{\mathbb{Q}}C^{\infty}(\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}}))\right]^{\operatorname{SO}_{\mathcal{V}}(\mathcal{k})\times K_{f}}\longrightarrow\Omega^{r}(Y)$ (3.32)
sending $\eta\otimes f$ to $\sum_{I}f(h_{I})\eta$ is an isomorphism, where $C^{\infty}(\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}}))$ is the space of locally constant functions. We define the homology and cohomology of $Y$ by
@@
sending a function $f$ to the form $\sum_{I\in\mathcal{C}}f(\cdot,h_{I})\varrho$. When $\eta$ is a top degree form, we can consider the integral over $Y$. Suppose that the form $\eta\in\Omega^{\dim(Y)}(Y)$ correspond to a function $f$ in the isomorphism (3.36). Then
@@ 3.3.  Kudla-Millson theta lift
Let $(p,q)$ be the signature of $\mathcal{V}_{\mathbb{R}}=\bigoplus\mathcal{V}_{\sigma}$. In [KM86, KM87] Kudla and Millson define a form
@@
Let us now define the Kudla-Millson theta series. Let $\varphi_{f}\in\mathcal{S}(\mathcal{V}_{\widehat{\mathcal{k}}})$ be the same Schwartz function as above, and let $K^{\prime}_{f}\subset\operatorname{SL}_{2}(\widehat{\mathcal{k}})$ be an open compact satisfying $\omega(k)\varphi_{f}(v)=\varphi_{f}(v)$ for every $k\in K^{\prime}_{f}$. Hence, the Schwartz function $\varphi_{f}$ is $K_{f}\times K^{\prime}_{f}$ invariant by the Weil representation. We define
@@
For $g\in\operatorname{SL}_{2}(\mathbb{A}_{\mathcal{k}})$ and $h_{f}\in\operatorname{SO}_{\mathcal{V}}(\widehat{\mathcal{k}})$, the Kudla-Millson theta series is
$\Theta_{\varphi}(g,h_{f})\coloneqq\sum_{v\in\mathcal{V}}(\omega(g,h_{f})\varphi)(v)\in C^{\infty}\left(\operatorname{SL}_{2}(\mathcal{k})\backslash\operatorname{SL}_{2}(\mathbb{A}_{\mathcal{k}})\right)\otimes\Omega^{q}(Y).$ (3.46)
@@
as a differential form on $Y$. Let $\Gamma^{\prime}\coloneqq\operatorname{SL}_{2}(\mathcal{k})\cap K^{\prime}_{f}$. For a point $(\tau_{1},\dots,\tau_{k})\in\mathscr{H}^{k}$ let $g_{\infty}=(g_{\tau_{1}},\dots,g_{\tau_{k}})\in\operatorname{SL}_{2}(\mathbb{R})^{k}$ where
@@
$\Theta_{\varphi}(\tau_{1},\dots,\tau_{k})$ $\coloneqq(y_{1}\cdots y_{k})^{-\frac{m}{2}}\Theta_{\varphi}(g_{\infty},1)\in\Omega^{q}(Y),$ (3.51)
where we recall that $m=\dim(\mathcal{V})$. By the work of Kudla and Millson - building on Weil’s construction of automorphic forms - the form $\Theta_{\varphi}$ transforms like a Hilbert modular form of parallel weight $m$ and level $\Gamma^{\prime}$, in the variables $\tau_{1},\dots,\tau_{k}$. They also show that the form is holomorphic in cohomology, in the sense that for every $\sigma$ we have
$\frac{\partial}{\partial\overline{\tau_{\sigma}}}\Theta_{\varphi}(\tau_{1},\dots,\tau_{k})=d\eta_{\sigma}$ (3.52)
@@
$\Theta_{\varphi}\in H^{q}(Y)\otimes M_{(m,\cdots,m)}(\Gamma^{\prime})$ (3.53)
where $M_{(m,\cdots,m)}(\Gamma^{\prime})$ is the space of Hilbert modular forms of parallel weight $m$ and level $\Gamma^{\prime}$. In particular, if $C\in\mathcal{Z}_{q}(Y)$ is a cycle, then
@@
$=\sum_{I\in\mathcal{C}}\int_{C_{I}}\Theta_{\varphi}(\tau_{1},\dots,\tau_{k},h_{I})\in M_{(m,\cdots,m)}(\Gamma^{\prime}).$ (3.54)
@@
$\Theta_{\varphi}(\tau_{1},\dots,\tau_{k},C)=\int_{Y}\Theta_{\varphi}(\tau_{1},\dots,\tau_{k})\wedge\eta_{C}\in M_{(m,\cdots,m)}(\Gamma^{\prime}).$ (3.55)
3.3.2.
Let us compute the level $\Gamma^{\prime}$, in the case $\mathcal{k}=\mathbb{Q}$. Let $L\subset V$ be an even integral lattice, in the sense that $Q(L,L)\subset 2\mathbb{Z}$. Consider the dual lattice
$L^{\vee}\coloneqq\left\{\left.v\in V\ \right|Q(v,w)\in\mathbb{Z}\ \textrm{for every }w\in L\right\}.$ (3.56)
Note that $L$ is a finite index subgroup of $L^{\vee}$, and let $N\coloneqq[L^{\vee}\colon L]$ be the level of $L$.
Let $\varphi_{f}=\bigotimes_{p}\varphi_{p}\in\mathcal{S}(V_{\widehat{\mathbb{Q}}})$ be the finite Schwartz function where $\varphi_{p}=\mathbf{1}_{L_{p}}\in\mathcal{S}(V_{\mathbb{Q}_{p}})$ is the characteristic function of the lattice $L_{p}=L\otimes_{\mathbb{Z}}\mathbb{Z}_{p}\subset V_{\mathbb{Q}_{p}}$. Note that $L=V_{\mathbb{Q}}\cap\widehat{L}$ where $\widehat{L}=\prod_{p}L_{p}$. We can define the dual lattice $L_{p}^{\vee}$ and the level $N_{p}\coloneqq[L_{p}^{\vee}\colon L_{p}]$ similarly. Note that $N_{p}=1$ for almost all primes $p$ and that $N=\prod_{p}N_{p}$.
Proposition 3.1.
The modular form $\Theta_{\varphi}(\tau,C)$ is of level $\Gamma_{0}(N)$.
Proof.
We have $\Gamma_{0}(N)=\operatorname{GL}_{2}(\mathbb{Q})^{+}\cap K_{0}(N)$ where
$K_{0}(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{GL}_{2}(\widehat{\mathbb{Z}}),N\mid c\right\}.$ (3.59)
Thus, we have to show that $\varphi_{p}$ is invariant by $K_{0}(N)_{p}=K_{0}(N_{p})_{p}\subset\operatorname{GL}_{2}(\mathbb{Z}_{p})$ under the Weil representation. The group $\operatorname{GL}_{2}(\mathbb{Z}_{p})$ is generated by matrices of the form
$S=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\ \begin{pmatrix}\mathbb{Z}_{p}^{\times}&0\\ 0&\mathbb{Z}_{p}^{\times}\end{pmatrix},\ \begin{pmatrix}1&\mathbb{Z}_{p}\\ 0&1\end{pmatrix}.$ (3.66)
Note that the action of $S$ is
$\omega(S)\varphi_{p}=\operatorname{vol}(L_{p})\mathbf{1}_{L_{p}^{\vee}}.$ (3.67)
In particular, if $N_{p}=1$, we have $L_{p}=L_{p}^{\vee}$ and $\operatorname{vol}(L_{p})=\operatorname{vol}(L_{p}^{\vee})^{-1}=\operatorname{vol}(L_{p})^{-1}=1$. Hence, the Schwartz function is preserved by $S$ when $N_{p}=1$. It is immediate to check that it is also invariant under the diagonal and unipotent matrices in (3.66), hence $\varphi_{p}$ is preserved by $K_{0}(N)_{p}=\operatorname{GL}_{2}(\mathbb{Z}_{p})$ when $N_{p}=1$. Suppose $N_{p}=p^{r}$ for some $r>0$ and write a matrix in $K_{0}(N)_{p}$ as
$\begin{pmatrix}a&b\\ p^{r}c&d\end{pmatrix}=S^{-1}\begin{pmatrix}1&\frac{p^{r}c}{a}\\ 0&1\end{pmatrix}S\begin{pmatrix}a&b\\ 0&d-\frac{bcp^{r}}{a}\end{pmatrix}.$ (3.74)
Note that $a\in\mathbb{Z}_{p}^{\times}$. Let us denote by $g_{1}$ and $g_{2}$ the two matrices in $\operatorname{GL}_{2}(\mathbb{Z}_{p})$ such that the right handside of (3.74) is $S^{-1}g_{1}Sg_{2}$. A direct computation shows that $\varphi_{p}$ is invariant by $g_{2}$, and that $\omega(S)\varphi_{p}=\operatorname{vol}(L_{p})\mathbf{1}_{L_{p}^{\vee}}$ is invariant under $g_{1}$. Hence we have $\omega(S^{-1}g_{1}Sg_{2})\varphi_{p}=\varphi_{p}$. ∎

### v4

@@ Acknowledgments.
for $x\in E$. Explicitely we have $\epsilon_{Q}(x)=A^{-1}_{Q}x^{T}A_{Q}$ where $A_{Q}$ denotes the Gram matrix of $Q$ and $x^{T}$ is the transpose of $x$. Let $E_{T}\subset\operatorname{End}(V)$ be the subalgebra consisting of all the $\mathbb{Q}$-endomorphisms $x\in\operatorname{End}(V)$ such that $xt=tx$ for any $t\in{\operatorname{T}}(\mathbb{Q})$. We will denote it simply by $E$ since there is no risk of confusion, and let $F$ be the subalgebra fixed by $\epsilon_{Q}$. We view $\operatorname{T}(\mathbb{Q})\subset\operatorname{SO}_{V}(\mathbb{Q})\subset\operatorname{End}(V)$. The following proposition is proved in [7, Proposition. 3.3].
@@
Over $\overline{\mathbb{Q}}$, the quadratic space $V$ is isomorphic to $\overline{\mathbb{Q}}^{2d}$ with quadratic form $Q(v,v)=v_{1}v_{d+1}+v_{2}v_{d+2}+\cdots+v_{d}v_{2d}$ i.e. with Gram matrix $\begin{pmatrix}0&\mathbf{1}_{d}\\ \mathbf{1}_{d}&0\end{pmatrix}$. With respect to this quadratic form, the torus can be diagonalized in $\operatorname{SO}_{V}(\overline{\mathbb{Q}})$ to
@@ 3.1. Weil representation
    the field $\mathcal{k}=F$ is an arbitrary totally real field of degree $k=d$. The quadratic space is $\mathcal{V}=E$, where $E$ is an étale algebra viewed as quadratic space of dimension $2m=2$ over $F$, and equipped with the quadratic form $Q_{\alpha}$. The orthogonal group $\operatorname{SO}_{\mathcal{V}}=\operatorname{SO}_{E}\simeq E^{1}$ is a torus.
@@
Let us describe this action more precisely, on a Schwartz function $\varphi=\varphi_{\infty}\otimes\varphi_{f}\in\mathcal{S}(\mathcal{V}_{\mathbb{R}})\otimes\mathcal{S}(\mathcal{V}_{\mathbb{A}_{f}})$. For $h\in\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k}})$, we have
@@
$(\omega(g,1)\varphi)(v)$ $={\chi_{w}\left(\frac{b\mathcal{Q}(v,v)}{2}\right)}\varphi(v).$ (3.8)
@@
For every real place $\sigma$ of $\mathcal{k}$, let $\mathcal{V}_{\sigma}\coloneqq\mathcal{V}\otimes_{\sigma}\mathbb{R}$ and $\mathcal{V}_{\mathbb{R}}\coloneqq\bigoplus\mathcal{V}_{\sigma}$. It is a real quadratic space, and let $(p_{\sigma},q_{\sigma})$ be its signature. Let $\mathcal{D}_{\sigma}^{+}$ be the associated connected symmetric space, that can be described as one of the two connected components of the Grassmanian $\mathcal{D}_{\sigma}$ of $q_{\sigma}$-dimensional oriented subspaces $z\subset\mathcal{V}_{\sigma}$ that are negative, i.e. ${\left.\kern-1.2pt\mathcal{Q}\vphantom{\big|}\right|_{z}}<0$. We set $\mathcal{D}^{+}=\prod\mathcal{D}_{\sigma}^{+}$, where the product ranges over the archimedean places of $\mathcal{k}$ for which $p_{\sigma}q_{\sigma}$ is nonzero [^2] In the case where $p_{\sigma}=0$ or $q_{\sigma}=0$, the manifold $\mathcal{D}_{\sigma}^{+}$ is just a point.. The dimension of $\mathcal{D}^{+}_{\sigma}$ is $p_{\sigma}q_{\sigma}$, so that the dimension of $\mathcal{D}^{+}$ is $\sum_{\sigma}p_{\sigma}q_{\sigma}$. At every place $\sigma$ let $\mathcal{k}_{\sigma}$ be the completion of $\mathcal{k}$, so that $\mathcal{k}_{\infty}\simeq\prod_{\sigma}\mathcal{k}_{\sigma}$. Since we assumed that $\mathcal{k}$ only has real places, we have $\mathcal{k}_{\sigma}\simeq\mathbb{R}$. At the level of the orthogonal group, we have $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\infty})\simeq{\prod_{\sigma}\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\sigma})}$, where $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\sigma})\simeq\operatorname{SO}(p_{\sigma},q_{\sigma})$. The connected component $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\infty})^{+}\simeq\prod_{\sigma}\operatorname{SO}(p_{\sigma},q_{\sigma})^{+}$ of the identity acts transitively on $\mathcal{D}^{+}$. Let $K_{\infty}$ be a maximal connected compact subgroup of $\operatorname{SO}_{\mathcal{V}}(\mathcal{k}_{\infty})^{+}$, that is isomorphic to $\prod_{\sigma}\operatorname{SO}(p_{\sigma})\times\operatorname{SO}(q_{\sigma})$. Hence
@@
is the space mentioned in the introduction. It has two connected components (corresponding to the two choices of orientations) and $\mathcal{D}^{+}$ is one of them.
@@
Let $\varphi_{f}\in\mathcal{S}(\mathcal{V}_{\mathbb{A}_{\mathcal{k},f}})$ be a finite Schwartz function, let $K_{f}\subset\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f})$ be an open compact preserving $\varphi_{f}$ in the sense that $\omega(k,1)\varphi_{f}=\varphi_{f}$. Let $K\coloneqq K_{\infty}K_{f}$ and consider the adelic space
$Y=\operatorname{SO}_{\mathcal{V}}(\mathcal{k})\backslash\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k}})/K\simeq\operatorname{SO}_{\mathcal{V}}(\mathcal{k})\backslash\mathcal{D}\times\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f})/K_{f}.$ (3.13)
@@
[…]
@@
$\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f})=\bigsqcup_{I\in\mathcal{C}}\operatorname{SO}_{\mathcal{V}}(\mathcal{k})^{+}h_{I}K_{f}.$ (3.15)
where $h_{I}\in\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f})$ are representants. The adelic space is a finite union of locally symmetric spaces
@@
Let $C^{\infty}(\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f}))$ be the space of locally constant functions on $\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f})$. The map
$\left[\Omega^{r}(\mathcal{D}^{+})\otimes_{\mathbb{Q}}C^{\infty}(\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f}))\right]^{\operatorname{SO}_{\mathcal{V}}(\mathcal{k})\times K_{f}}\longrightarrow\Omega^{r}(Y)$ (3.18)
sending $\eta\otimes f$ to $\sum_{I}f(h_{I})\eta$ is an isomorphism, where $C^{\infty}(\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f}))$ is the space of locally constant functions. We define the homology and cohomology of $Y$ by
@@
sending a function $f$ to the form $\sum_{I\in\mathcal{C}}f(\cdot,h_{I})\varrho$. When $\eta$ is a top degree form, we can consider the integral over $Y$. Suppose that the form $\eta\in\Omega^{\dim(Y)}(Y)$ corresponds to a function $f$ in the isomorphism (3.22). Then
@@ 3.3. Kudla-Millson theta lift
Let $(p,q)=\sum_{\sigma}(p_{\sigma},q_{\sigma})$} be the signature of $\mathcal{V}_{\mathbb{R}}=\bigoplus\mathcal{V}_{\sigma}$. In [14, 15] Kudla and Millson define a form
@@
Let us now define the Kudla-Millson theta series. Let $\varphi_{f}\in\mathcal{S}(\mathcal{V}_{\mathbb{A}_{\mathcal{k},f}})$ be the same Schwartz function as above, and let $K^{\prime}_{f}\subset\operatorname{SL}_{2}(\mathbb{A}_{\mathcal{k},f})$ be an open compact satisfying $\omega(1,k^{\prime})\varphi_{f}(v)=\varphi_{f}(v)$ for every $k^{\prime}\in K^{\prime}_{f}$. Hence, the Schwartz function $\varphi_{f}$ is $K_{f}\times K^{\prime}_{f}$ invariant by the Weil representation. We define
@@
For $g\in\operatorname{SL}_{2}(\mathbb{A}_{\mathcal{k}})$ and $h_{f}\in\operatorname{SO}_{\mathcal{V}}(\mathbb{A}_{\mathcal{k},f})$, the Kudla-Millson theta series is
[…]
@@
as a differential form on $Y$. Let $\Gamma^{\prime}_{\mathcal{k}}\coloneqq\operatorname{SL}_{2}(\mathcal{k})\cap K^{\prime}_{f}$. For a point $(\tau_{1},\dots,\tau_{k})\in\mathscr{H}^{k}$ let $g_{\infty}=(g_{\tau_{1}},\dots,g_{\tau_{k}})\in\operatorname{SL}_{2}(\mathbb{R})^{k}$ where
@@
$\Theta_{\varphi}(\tau_{1},\dots,\tau_{k})$ $\coloneqq(y_{1}\cdots y_{k})^{-\frac{m}{2}}\Theta_{\varphi}(g_{\infty},\cdot)\in\Omega^{q}(Y),$ (3.33)
where we recall that $m=\dim(\mathcal{V})$. By the work of Kudla and Millson - building on Weil’s construction of automorphic forms - the form $\Theta_{\varphi}$ transforms like a Hilbert modular form of parallel weight $m$ and level $\Gamma^{\prime}_{\mathcal{k}}\subset\operatorname{SL}_{2}(\mathcal{k})$, in the variables $\tau_{1},\dots,\tau_{k}$. They also show that the form is holomorphic in cohomology, in the sense that for every $\sigma$ we have
$\frac{\partial}{\partial\bar{\tau}_{\sigma}}\Theta_{\varphi}(\tau_{1},\dots,\tau_{k})=d\eta_{\sigma}$ (3.34)
@@
$\Theta_{\varphi}\in H^{q}(Y)\otimes M_{(m,\cdots,m)}(\Gamma^{\prime}_{\mathcal{k}})$ (3.35)
where $M_{(m,\cdots,m)}(\Gamma^{\prime}_{\mathcal{k}})$ is the space of Hilbert modular forms of parallel weight $m$ and level $\Gamma^{\prime}_{\mathcal{k}}\subset\operatorname{SL}_{2}(\mathcal{k})$. In particular, if $C\in\mathcal{Z}_{q}(Y)$ is a cycle, then
@@
$=\sum_{I\in\mathcal{C}}\int_{C_{I}}\Theta_{\varphi}(\tau_{1},\dots,\tau_{k},h_{I})\in M_{(m,\cdots,m)}(\Gamma^{\prime}_{\mathcal{k}}).$ (3.36)
@@
$\Theta_{\varphi}(\tau_{1},\dots,\tau_{k},C)=\int_{Y}\Theta_{\varphi}(\tau_{1},\dots,\tau_{k})\wedge\eta_{C}\in M_{(m,\cdots,m)}(\Gamma^{\prime}_{\mathcal{k}}).$ (3.37)